1.A firm has a capital structure that is 30% Debt, and 70% equity. Its YTM on current bonds is 9%, and its tax rate is 25%. If the WACC is 10.5%, what is the cost of equity?

2.A firm has a 30% tax rate, and a 3% cost to issue new common stock.

It has determined the optimal capital structure as shown in the table below. What is the firm's cost of capital?

Type of Capital Capital Structure Cost of Capital WACC
$000s % of total capital
Debt 15000 0.06
Retained Earnings 47000 0.12
New Common Stock 22000
Total Capital
Answer #1

Hi

As per policy we will solve only first question here.

here D/V = 0.3

E/V =0.7

WACC = (D/V)*(1-t)*cost of debt + E/V*cost of equity

10.5 = 0.3*(1-0.25)*9 + 0.7*cost of equity

10.5 = 2.025 +0.7*cost of equity

cost of equity = 8.475/0.7 = 12.11%

Thanks